OpenStudy (anonymous):

One model for home appreciation is as follows, where C is the original cost, V is the value of the home after t years, and b is the rate of appreciation. V = Ce^bt If the original cost of the home was $185,000 and the rate of appreciation is 8%, what will be the value of the home after 8 years? [Round to the nearest dollar.] $350,849 $197,227 $1,184,000 $200,408

OpenStudy (anonymous):

Is e a part of the formula?

OpenStudy (anonymous):

ya i believe so

OpenStudy (anonymous):

If so what do it represent? They have mentioned nothing about e in the description right?

OpenStudy (anonymous):

Ha there u go

OpenStudy (anonymous):

It is Ce^bt earlier u mentioned it Cebt which confused me :)

OpenStudy (anonymous):

So now we have C as $185,000 b as 8% which can be 8/100 and t as 8 years V = 185000*e^(8/100*8) V = 185000*e^(64/100) = 185000*e^(0.64) Simplify it and you get the answer

OpenStudy (anonymous):

A